ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60G35 Khalil*, Z., and M. E. Ghitany**, * Concordia Univ., Montreal. ,** Kuwait Univ., Kuwait On the total amount of lost information in a TASI system Consider the process of information transmission in a Time Assignment Speech Interpolation (TASI) system . The functioning of such systems is usually characterized by the fraction of freeze-out talkspurts of talkers . In the present article we analyze the behavior of such systems and derive characteristics for the total amount of lost information in a TASI system, under the assumption that input subscribers are continuously occupied, talkspurts and silence periods are exponentially distributed.